# Brilla AI: AI Contestant for the National Science and Maths Quiz

### v1

@@ 8 Conclusion
In this work, we built Brilla AI, an AI contestant that we deployed to unofficially compete remotely and live in the Riddles round of the 2023 NSMQ Grand Finale, the first of its kind in the 30-year history of the competition. Our AI answered one of the 4 riddles ahead of the 3 human contesting teams, unofficially placing second (tied). Our next step is to improve the performance of our system ahead of 2024 NSMQ during which we plan to unofficially compete in both Round 5 (Riddles) and Round 4 (True or False) for all 5 rounds of the competition. Aside from being an interesting intellectual challenge, we are building an AI that addresses the unique context of Africans — transcribes speech with a Ghanaian accent, provides answers to scientific questions drawn from a Ghanaian quiz, and says answers with a Ghanaian accent — that could be integrated into an education tool. Imagine a student in a rural part of Ghana calling a toll-free number with a feature phone, asking this AI numerous science questions, and the AI understands her even with her Ghanaian accent and then provides explanations using local context in a Ghanaian accent. That vision is our long-term goal and would radically transform learning support for students across Africa, enabling millions of young people to have one-on-one interactions and support even with limited access to teachers, computers, or even smartphones leading to the democratizing of science education across Africa!

### v2

@@ 8 Conclusion
In this work, we built Brilla AI, an AI contestant that we deployed to unofficially compete remotely and live in the Riddles round of the 2023 NSMQ Grand Finale, the first of its kind in the 30-year history of the competition. Our AI answered one of the 4 riddles ahead of the 3 human contesting teams, unofficially placing second (tied). Our next step is to improve the performance of our system ahead of 2024 NSMQ during which we plan to unofficially compete in both Round 5 (Riddles) and Round 4 (True or False) for all 5 stages of the competition. Aside from being an interesting intellectual challenge, we are building an AI that addresses the unique context of Africans — transcribes speech with a Ghanaian accent, provides answers to scientific questions drawn from a Ghanaian quiz, and says answers with a Ghanaian accent — that could be integrated into an education tool. Imagine a student in a rural part of Ghana calling a toll-free number with a feature phone, asking this AI numerous science questions, and the AI understands her even with her Ghanaian accent and then provides explanations using local context in a Ghanaian accent. That vision is our long-term goal and would radically transform learning support for students across Africa, enabling millions of young people to have one-on-one interactions and support even with limited access to teachers, computers, or even smartphones leading to the democratizing of science education across Africa!